Cards are drawn, one at a time, from a standard deck; each card is replaced before the next one is drawn. Let Y be the number of draws necessary so that ten of the draws result in spades. Find E(Y).

The answer is 40, I'm just wondering how to get it.
Answer #1

here this is negative binomial distribution with parameter p=probability of spade =1/4 and r =number of success =10

therefore expected number of cards to be drawn =r/p =10/(1/4) =40